If the cofactor matrix is



7 15 9



6 74



10 2 7



−



 



 



−



 



 



−



 



, then the Adjoint matrix is


1
Expert's answer
2021-11-25T14:17:29-0500

The adjoint matrix is the transpose of the cofactor matrix.


"\begin{bmatrix}
 7 & 15 & 9 \\
 6 & 7 & 4 \\
 10 & 2 & 7
\end{bmatrix}^T=\begin{bmatrix}
 7 & 6 & 10 \\
 15 & 7 & 2 \\
 9 & 4 & 7
\end{bmatrix}"

Answer


"\begin{bmatrix}
 7 & 6 & 10 \\
 15 & 7 & 2 \\
 9 & 4 & 7
\end{bmatrix}"